# Investigating Guiding Information for Adaptive Collocation Point Sampling in PINNs

### v1

@@ 1.2 PINNs
It is important to distribute these points throughout the domain however evidence suggests that it may also be beneficial to identify and place a greater proportion of points in particular regions of importance, so that the loss term especially includes these areas. This can ensure that optimising the cost function improves how well the PDE is satisfied around these locations, and therefore hopefully improves the solution in those areas.
@@ 2.1 Probability density functions
Here $Y(\mathcal{X})$ is the vector containing the magnitude of the information source chosen (e.g. the current residual or its derivatives, or some derivatives of the current estimate of the solution), $\overline{Y(\mathcal{X})^{k}}$ is the mean value across all points in $\mathcal{X}$ and $k,c$ are constant hyper-parameters that affect the resampling behaviour. A visualisation of the distribution of points obtained using the PDE residual as $Y(\mathcal{X})$ for solving the Burgers’ Equation is shown in figure 1
@@ 2.2 Problem Definition
Figure 2: Burgers’ Equation contour plot of example solution with $\nu=\frac{\pi}{100}$
@@ 2.3 Results: default case
Figure 3: Error against number of resamples solving Burgers’ Equation with described default parameters $v=\frac{\pi}{100};N=2000$
The initial distribution is seen to have an effect in both the fixed and residual-based methods, with the use of the Hammersley sampling algorithm resulting in consistently lower error. The local residual method with random initialisation ($PDE,R$) exhibited the least robustness and greatest variability, which can be seen from the fact increasing the number of resamples did not consistently lower the average error between runs. Generally, more resamples does result in improving accuracy, with error descending below 0.1% for above 30 resamples for the most of the adaptive resampling methods.
For the same case we also look at the error achieved using different numbers of collocation points. This was obtained for 100 rounds of resampling and is plotted in figure 4. This figure suggests that the threshold for the number of points required to obtain results of a given accuracy varies from method to method. Even for the most accurate method ($PDE_{xt}$), the error seems to plateau however between 0.01% and 0.1%, which the other adaptive methods eventually reach at the default 2000 points. Nevertheless, this clearly suggests that the $PDE_{xt}$ approach could allow similar accuracy to other techniques but at a significantly lower cost.
Figure 4: Error against number of collocation points for different resampling methods solving Burgers’ Equation with described default parameters $v=\frac{\pi}{100};$ 100 resamples
@@ 3.2 Adjusting PDE parameters
For the following example, we investigate adjusting the value of the $\nu$ parameter in the PDE; which affects the magnitude of the diffusion term $\nu\;u_{xx}$. Lower values increase the shock sharpness, increasing the complexity of the problem without fundamentally changing the shape of the solution. Increasing $\nu$ sees the opposite effect. We use values of $\nu$ of $0.01$ and $0.001$, approximately multiplying and dividing by 3 from the default $\nu=\frac{\pi}{100}$ used in the previous section. The ground truth is again computed using high-resolution numerical simulations and the error when using different information sources to control the adaptive resampling is again compared against the fixed uniform methods: see figures 6 and 7.

### v2

@@ 1.2 PINNs
It is important to distribute these points throughout the domain, however, evidence suggests that it may also be beneficial to identify and place a greater proportion of points in particular regions of importance, so that the loss term especially includes these areas. This can ensure that optimising the cost function improves how well the PDE is satisfied around these locations, and therefore hopefully improves the solution in those areas.
@@ 2.1 Probability density functions
Here $Y(\mathcal{X})$ is the vector containing the magnitude of the information source chosen (e.g. the current residual or its derivatives, or some derivatives of the current estimate of the solution), $\overline{Y(\mathcal{X})^{k}}$ is the mean value across all points in $\mathcal{X}$ and $k,c$ are constant hyper-parameters that affect the resampling behaviour. A visualisation of the distribution of points obtained using the PDE residual as $Y(\mathcal{X})$ for solving the Burgers’ Equation is shown in figure 1.
@@ 2.2 Problem Definition
Figure 2: Burgers’ Equation contour plot of example solution with $\nu=\frac{1}{100\pi}$
@@ 2.3 Results: default case
Figure 3: Error against number of resamples solving Burgers’ Equation with described default parameters $v=\frac{1}{100\pi};N=2000$
The initial distribution is seen to have an effect in both the fixed and residual-based methods, with the use of the Hammersley sampling algorithm resulting in consistently lower error. The local residual method with random initialisation ($PDE,R$) exhibited the least robustness and greatest variability, which can be seen from the fact increasing the number of resamples did not consistently lower the average error between runs. Generally, more resamples does result in improving accuracy, with error descending below 0.1% for above 30 resamples for most of the adaptive resampling methods.
For the same case, we also look at the error achieved using different numbers of collocation points. This was obtained for 100 rounds of resampling and is plotted in figure 4. This figure suggests that the threshold for the number of points required to obtain results of a given accuracy varies from method to method. Even for the most accurate method ($PDE_{xt}$), the error seems to plateau between 0.01% and 0.1%, which the other adaptive methods eventually reach at the default 2000 points. Nevertheless, this clearly suggests that the $PDE_{xt}$ approach could allow similar accuracy to other techniques but at a significantly lower cost.
Figure 4: Error against number of collocation points for different resampling methods solving Burgers’ Equation with described default parameters $v=\frac{1}{100\pi};$ 100 resamples
@@ 3.2 Adjusting PDE parameters
For the following example, we investigate adjusting the value of the $\nu$ parameter in the PDE; which affects the magnitude of the diffusion term $\nu\;u_{xx}$. Lower values increase the shock sharpness, increasing the complexity of the problem without fundamentally changing the shape of the solution. Increasing $\nu$ sees the opposite effect. We use values of $\nu$ of $0.01$ and $0.001$, approximately multiplying and dividing by 3 from the default $\nu=\frac{1}{100\pi}$ used in the previous section. The ground truth is again computed using high-resolution numerical simulations and the error when using different information sources to control the adaptive resampling is again compared against the fixed uniform methods: see figures 6 and 7.